OpenStudy (anonymous):

Find the slope of the following curve at x=5 y= 1/(x-2) No idea how to do really.explaination needed

OpenStudy (anonymous):

slope is the value of first derivative at the indicated point

OpenStudy (anonymous):

I thought that..just no idea how to find here. havent done it yet

OpenStudy (anonymous):

do u know how to find the derivative of a function?

OpenStudy (anonymous):

y=(x-2)^(-1) y'=(-1)(x-2)^(-2) just put x=5 to find slope

OpenStudy (anonymous):

What? we were told to do like. F(xo+h)-F(x)/h

OpenStudy (anonymous):

yea right, but that is by definition.here i used power rule to find the derivative

OpenStudy (anonymous):

if y=x^n y'=nx^(n-1)

OpenStudy (anonymous):

No idea how that works..but i guess it does. Thanks Ill have to ask professor. I dont get how you got the y1=(-1)(x-2)^(-2)

OpenStudy (anonymous):

ok, i mentioned the power rule that i used,